OpenStudy (anonymous):

immy is writing a paper for one of his classes. The paper has to be 2,500 words long, and so far he has written 700 words. If he only has 5 more days to write his paper and wants to write the same number of words each day, then how many words must he write per day to finish the paper? 360 860 390 1,800

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

360

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

John has 12 boxes of apples. Each box holds 6 apples. If 5 of the boxes are full, and 7 of the boxes are half full, how many apples does John have? 36 51 72 24

OpenStudy (anonymous):

51

OpenStudy (texaschic101):

5 boxes are full.....5 * 6 = 30 apples 7 boxes are half full....7 * 3 = 21 30 + 21 = 51

OpenStudy (anonymous):

Angie read 2 books in 5 days, and each book contained 350 pages. If Angie read the same number of pages each day, how many pages did she read per day? 77 350 700 140

OpenStudy (texaschic101):

2 books with 350 pgs = 700 pgs 700/5 = 140 pgs per day

OpenStudy (anonymous):

Lisa is on a trip of 2,897 miles. She has already traveled 90 miles. She has 7 days left in her trip. How many miles does she need to travel each day to complete the trip? 2,807 491 401 2,406

OpenStudy (anonymous):

2897-90=2807 2807/7=401

OpenStudy (texaschic101):

you beat me to that one...lol

OpenStudy (anonymous):

lol, i have a stat question if you can help

OpenStudy (texaschic101):

oh...not real good with those

OpenStudy (anonymous):

go

OpenStudy (anonymous):

me either

OpenStudy (anonymous):

u guys rock

OpenStudy (anonymous):

but i am going to bed

OpenStudy (texaschic101):

ok...good night :)

OpenStudy (anonymous):

texaschic101 can u metal me

OpenStudy (texaschic101):

not a problem :)

OpenStudy (texaschic101):

sweet dreams :)